Finest funiture sell consumers enough furniture for a whole house in 1 single transaction. Finest thinks the company needs a larger space for storage. There are 2 possible locations.

Location Y has annual fixed costs of $115,000 and variable costs of $9,000 per furniture set; location Z has annual fixed costs of $315,000 and variable costs of $5,000 per set. The sets sell for $14,000.


a. at what output level would finest be indifferent between the 2 potential spots?


  Output level    units


b-1.

at what output range would Y be the prefernce? (Enter your answer as a whole number. Do not include the indifference point in your answer.)


Output range 0 to   


b-2.

At what output range would Z be the preference? (Enter your answer as a whole number. Do not include the indifference point in your answer.)


  Output range or more
Answer #1

for indifference ;

profit for Y =profit for Z

(14000-9000)*x-115000=(14000-5000)*x-315000

4000x=20000

Output level x=50    units

b-1)

Output range =0 to 50 units

b-2)   Output range =50 or more